Consider the following regression results based on 20 observations. [You may find it useful to reference the t table.]

Coefficients Standard Error t Stat p-value
Intercept 31.4171 4.4496 7.061 0.000
x1 0.1629 0.1944 0.838 0.413

a-1. Choose the hypotheses to determine if the intercept differs from zero.

  • H0: β0 ≥ 0; HA: β0 < 0

  • H0: β0 = 0; HA: β0 ≠ 0

  • H0: β0 ≤ 0; HA: β0 > 0


a-2. At the 5% significance level, what is the conclusion to the hypothesis test? Does the intercept differ from zero?

  • Reject H0; the intercept is greater than zero.

  • Reject H0; the intercept differs from zero.

  • Do not reject H0; the intercept is greater than zero.

  • Do not reject H0; the intercept differs from zero.

b-1. Construct the 95% confidence interval for the slope coefficient. (Negative values should be indicated by a minus sign. Round "tα/2,df" value to 3 decimal places and final answers to 2 decimal places.)

Confidence interval _____ to ______?


b-2. At the 5% significance level, can we conclude the slope differs from zero?

  • Yes, since the interval contains zero.

  • Yes, since the interval does not contain zero.

  • No, since the interval contains zero.

  • No, since the interval does not contain zero.
Answer #1

a-1. Null and alternative hypothesis:

H0: β0 = 0; HA: β0 ≠ 0

a-2. p-value = 0.000

Reject H0; the intercept differs from zero.

b-1. 95% confidence interval for the slope coefficient.

Critical value, t_c = T.INV.2T(0.05, 18) = 2.1009

95% Confidence interval for slope:

Lower limit = b1 - tc*se(b1) = 0.1629 - 2.1009*0.1944 = -0.25

Upper limit = b1 + tc*se(b1) = 0.1629 + 2.1009*0.1944 = 0.57

b-2. No, since the interval contains zero.